# Benchmarking 3D Human Pose Estimation Models under Occlusions

### v2

@@ 4.1.3 Occlusion Noise Analysis
Additionally, the results in Table 3 and Figure 5 reinforce this categorization. For $\sigma=\{0.001,0.005\}$, many models maintain or slightly improve their performance compared to using ground-truth 2D input. This suggests that such low levels of noise reflect the standard deviation typically present in training with detected 2D poses. At $\sigma=0.01$, the noise magnitude begins to simulate mild occlusion behavior while still partially overlapping with the visible range. Once the noise increases to $\sigma=0.03$, the performance of the models drops more noticeably, reflecting a realistic degree of occlusion-induced perturbation.
From this analysis, we conclude that zero-mean Gaussian noise with $\sigma=0.03$ provides a reliable simulation of real occlusions in 2D keypoints, and will be used in the subsequent per-keypoint occlusion evaluation (Protocol 2).
4.2 Per-Keypoint Occlusion Analysis
In the second part of our study, we focus on identifying which keypoints are more vulnerable to occlusion and whether this behavior persists when the same keypoints are visible. Building on the detector noise and occlusion severity analyses (Sections 4.1 and 4.1.3), we selected $\sigma=0.03$ as a representative occlusion noise level, simulating moderate to severe occlusions.
@@ 4.2 Per-Keypoint Occlusion Analysis
To isolate the impact of each keypoint, in each experiment, we added zero-mean Gaussian noise into one individual 2D keypoint, between frames 50 and 149 of the first 243-frame segment. This localized occlusion simulation allows us to evaluate per-keypoint error propagation in the 3D predictions. We refer to this procedure as Protocol 2, described in Section 3.3.2.
@@
Figure 7 displays the aggregated MPJPE results across all models. We distinguish the error on the occluded keypoint, the remaining visible keypoints, and the full-pose average. From Figure 7 we can see that a common pattern emerges: certain keypoints are inherently more difficult to estimate when occluded. Notably, the wrists (Lwrist, Rwrist), feet (RFoot, LFoot), Nose and Head show higher MPJPE under occlusion. These extremities are small, fast-moving, and often heavily occluded, making them particularly error-prone. In contrast, more central and structurally constrained keypoints, such as the hips, knees, shoulders, and elbows, exhibit lower errors, even under occlusion.
Figure 7: MPJPE per keypoint across occluded frames, where each keypoint is occluded independently using Gaussian noise ($\sigma=0.03$). We report (i) MPJPE of the occluded keypoint, (ii) MPJPE of visible keypoints, and (iii) average pose error. Values are averaged across all models.
@@
We extended this analysis to real occlusions in the BlendMimic3D dataset using the provided labels and added Gaussian noise ($\sigma=0.03$). In Figure 9, we show the MPJPE of each keypoint when visible vs. occluded within real occlusion scenarios. The same pattern persists: error increases with occlusion, and the most affected keypoints remain the same. This supports our conclusion that keypoints with higher degrees of freedom (knees, feet, wrists, and elbows) are consistently harder to estimate.
[ht] Refer to caption
Figure 9: MPJPE for each keypoint under real occlusion conditions in BlendMimic3D (BM3D). We compare errors when the keypoint is visible versus occluded across poses. These reflect the natural distribution of occlusions and keypoint visibility. Values are averaged across all models.
A detailed per-model breakdown of all three analyses, per-keypoint occlusion (Figure 7), comparison with fully visible pose (Figure 8), and real occlusion analysis (Figure 9), is provided in the supplementary material. These results include model-specific plots of occluded vs. visible keypoint errors, enabling a deeper understanding of how each architecture responds to occlusion.
@@ 5 Beyond Occlusions: Evaluating Additional Factors
[…]
Figure 10: Data distribution of subject orientation and distance to the camera in Human3.6M (a) and BlendMimic3D (b).
@@ 5.2 Camera-Related Sensitivity
6 Challenges and Limitations
Throughout this study, several challenges and limitations of current 3D HPE models have emerged.
Skeleton format mismatch is a key issue. Most 3D HPE models are trained using the Human3.6M keypoint format, whereas real-world 2D detectors often output COCO-format keypoints. This mismatch can introduce errors unless a careful alignment and mapping are performed between the two formats. For our benchmark, we manually aligned the BlendMimic3D dkeleton format to match Human3.6M using Blender software.
We also identified an evaluation bias related to hip keypoint centering. Many models normalize poses by subtracting the 3D coordinates of the hip, effectively zeroing the hip and removing the global trajectory during training. However, some models, including P-STMO, D3DP, FinePose, and DTF, explicitly reset the predicted hip to zero during evaluation, regardless of the actual output. This post-processing step can artificially reduce MPJPE and introduces inconsistency across models. To ensure fairness, we excluded the hip keypoint from evaluation for all models.
One of the most critical limitations is occlusion handling. Most state-of-the-art models suffer a noticeable drop in performance when keypoints are partially or fully occluded. In particular, models like D3DP, despite showing strong results under clean conditions, tend to degrade rapidly as occlusion levels increase. This indicates that robustness to occluded inputs remains an open challenge in 3D HPE.
We also found that model performance varies significantly across body parts. Keypoints with greater degrees of freedom, such as the hands, feet, and head, consistently resulted in higher errors, especially under occlusion. This points to intrinsic difficulties in modeling highly dynamic keypoints.
Camera-related distribution shifts pose another challenge. Our analysis showed that models trained solely on Human3.6M struggle when evaluated under camera distances or viewing angles that were not present during training, particularly at closer distances. This reveals a clear gap in the generalization ability of current approaches.
Taken together, these findings show that despite strong performance on benchmark datasets, current 3D HPE models are not yet robust for real-world applications, especially in scenarios involving occlusion, diverse camera setups, and dynamic poses.
@@ 7 Conclusion
In this study, we conducted a comprehensive benchmark of state-of-the-art 2D-to-3D human pose estimation models under realistic occlusion scenarios using the BlendMimic3D dataset. Our focus was to evaluate model robustness to noisy 2D inputs, a common challenge in real-world applications.
@@
Our evaluation revealed several key findings. All models, regardless of architecture, suffer performance degradation under occlusion. Diffusion-based models, despite their probabilistic formulation, proved sensitive to noisy inputs since their denoising processes rely heavily on the 2D input pose. Occlusion-aware models such as DTF showed improved robustness under severe occlusion but underperformed on clean inputs. Protocol 2 revealed that keypoints with higher degrees of freedom, such as wrists and feet, are consistently more error-prone, even when not occluded.
Beyond occlusions, we highlighted that camera-related distribution shifts between training (Human3.6M) and evaluation (BlendMimic3D) affect generalization, especially at closer distances and unseen viewpoints. We also identified evaluation inconsistencies, such as hip keypoint normalization and skeleton format mismatches, which we addressed to ensure fair comparison across models.
Despite advances in 3D HPE, our results show that current models are not yet robust enough for real-world environments. Future research should focus on developing occlusion-aware models that not only generalize across diverse viewing conditions and skeleton formats, but also maintain strong performance under clean input settings.
@@ Appendix C Beyond Occlusions: Action Variability (Per Model)
In addition to camera-related analysis, we also investigate the impact of action variability. Different actions include varying poses and spatial configurations, which may affect prediction accuracy. Figure 16 reports the MPJPE for each action across all camera viewpoints, averaged across all models, while Figure 17 presents the corresponding per model performances. These figures show that the test set includes a broad and realistic distribution of actions and viewpoints (see also Figure 10), supporting the diversity needed to evaluate generalization.

### v3

@@ 4.3 Protocol 2: Per-Keypoint Occlusion Analysis
Results in Table 4 and Figure 5 show that for $\sigma=\{0.001,0.005\}$, many models maintain or slightly improve their performance compared to using ground-truth 2D input. This suggests that such low levels of noise reflect the standard deviation typically present in training with detected 2D poses. At $\sigma=0.01$, the noise magnitude begins to simulate mild occlusion behavior while still partially overlapping with the visible range. Once the noise increases to $\sigma=0.03$, the performance of the models drops more noticeably, reflecting a realistic degree of occlusion-induced perturbation. Based on this, we select $\sigma=0.03$ as a representative occlusion noise level, simulating moderate to severe occlusions.
Figure 7: MPJPE per keypoint across occluded frames, where each point in the plot corresponds to the average (dashed line) and standard deviation of 10 runs where Gaussian noise ($\sigma=0.03$) is added solely to the corresponding keypoint.
@@
We now focus on identifying which keypoints are more vulnerable to occlusion and whether this behavior persists when the same keypoints are visible. So, to isolate the impact of each keypoint, in each experiment, we added zero-mean Gaussian noise ($\sigma=0.03$) into one individual 2D keypoint, between frames 50 and 149 of the first 243-frame segment. This localized occlusion simulation allows us to evaluate per-keypoint error propagation in the 3D predictions.
@@
Figure 7 displays the aggregated MPJPE results across all models. We can see that a common pattern emerges: certain keypoints are inherently more difficult to estimate when occluded. Notably, the wrists (Lwrist, Rwrist), feet (RFoot, LFoot), Nose and Head show higher MPJPE under occlusion. These extremities are small, fast-moving, and often heavily occluded, making them particularly error-prone. In contrast, more central and structurally constrained keypoints, such as the hips, knees, shoulders, and elbows, exhibit lower errors, even under occlusion.
@@
Figure 9: MPJPE for each keypoint under realistic occlusion conditions in BlendMimic3D (BM3D). We compare errors when the keypoint is visible versus occluded across poses. These reflect the natural distribution of occlusions and keypoint visibility. Values are averaged across all models.
We extended this analysis to realistic occlusion scenarios in BlendMimic3D using the provided labels and added Gaussian noise ($\sigma=0.03$). Figure 9 shows the MPJPE of each keypoint when visible vs. occluded. The same pattern persists: error increases with occlusion, and the most affected keypoints remain the same. This supports our conclusion that keypoints with higher degrees of freedom (knees, feet, wrists, and elbows) are consistently harder to estimate.
A detailed per-model breakdown of all three analyses, per-keypoint occlusion (Figure 7), comparison with fully visible pose (Figure 8), and real occlusion analysis (Figure 9), is provided in the supplementary material. These results include model-specific plots of occluded vs. visible keypoint errors, enabling a deeper understanding of how each architecture responds to occlusion. Importantly, all models exhibit the same qualitative pattern described here.
@@ 5 Beyond Occlusions: Evaluating Additional Factors
Figure 10: Data distribution of subject orientation and distance to the camera in Human3.6M (a) and BlendMimic3D (b). White pixels indicate no samples.
@@
To better understand how domain shift affects model generalization, we compare the distributions of Human3.6M and BlendMimic3D in terms of the subject’s orientation and distance relative to the camera. Using the 3D poses and camera parameters, we estimate the Euclidean distance between the subject’s center (mean of all 3D keypoints) and the camera position. The orientation is measured as the angle between the camera’s viewing direction (Z-axis) and the normal vector of the subject’s shoulder-hip plane, defined using the hip and both shoulder keypoints. An orientation of $0\text{\,}\mathrm{\SIUnitSymbolDegree}$ indicates the subject is facing away from the camera, while $180\text{\,}\mathrm{\SIUnitSymbolDegree}$ corresponds to facing directly toward the camera.
Figure 10 presents the data distribution of orientation and distance for training data from Human3.6M (subject S1, S5, S6, S7, and S8) and the test data from BlendMimic3D (subject S2). As shown, Human3.6M primarily captures subjects facing the camera, turning away, or viewed from the side, between 4 and 6 meters, with two dominant peaks around $20\text{\,}\mathrm{\SIUnitSymbolDegree}$ and $160\text{\,}\mathrm{\SIUnitSymbolDegree}$, highlighting limited diversity in viewpoint. In contrast, BlendMimic3D exhibit a broader and more uniform distribution across both orientation and distance. This makes it a comprehensive benchmark for evaluating model robustness. Notably, BlendMimic3D includes poses at closer distances, scenarios that were unseen during training.
@@ 6 Conclusion
In this study, we benchmarked SOTA 2D-to-3D HPE models under realistic occlusion scenarios using the BlendMimic3D dataset. Our focus was to evaluate model robustness to noisy 2D inputs, a common real-world challenge.
@@
Results showed that all models degrade under occlusion. Diffusion-based models proved sensitive to noisy inputs since their denoising processes rely heavily on the 2D input pose. This highlights the need for uncertainty-aware or visibility-conditioned denoising mechanisms, as well as sampling strategies that reduce dependence on potentially corrupted inputs. Occlusion-aware models such as DTF showed improved robustness under severe occlusion but underperformed on clean inputs. Overall, no model was able to maintain consistent performance across the full occlusion spectrum, however PoseFormerV2 and DTF exhibited the strongest resilience under moderate and severe corruption, respectively. This suggests that hybrid architectures combining frequency-domain smoothing with confidence-aware fusion should be explored.
Protocol 2 revealed that keypoints with higher degrees of freedom, such as wrists and feet, are consistently more error-prone, even when not occluded. Therefore, future work should incorporate explicit per-joint modeling, such as local corrective branches or learned keypoint-reliability modules.
Beyond occlusions, we observed that camera-related distribution shifts between training (Human3.6M) and evaluation (BlendMimic3D) affect generalization, especially at closer distances and unseen viewpoints. We also identified and addressed evaluation inconsistencies, such as hip keypoint normalization and skeleton format mismatches, to ensure fairness.
Despite advances in 3D HPE, our results show that current models are not yet robust enough for real-world environments. Future research should focus on occlusion-aware models that generalize across diverse viewing conditions and skeleton formats, while maintaining strong performance on clean inputs.
7 Acknowledgments
This work was supported by research grant 10.54499/2022.07849.CEECIND/CP1713/CT0001 and LARSyS funding (DOI: 10.54499/LA/P/0083/2020, 10.54499/UIDP/50009/2020 and 10.54499/UIDB/50009/2020), through Fundação para a Ciência e a Tecnologia (FCT), and by the SmartRetail project [PRR- C645440011-00000062], through IAPMEI- Agência para a Competitividade e Inovação. Filipa Lino is supported by the FCT doctoral grant [2025.03757.BD].
@@ Appendix C Beyond Occlusions: Action Variability (Per Model)
In addition to camera-related analysis, we also investigate the impact of action variability. Different actions include varying poses and spatial configurations, which may affect prediction accuracy. Figure 16 presents the per-model performance across actions, highlighting how each architecture responds to this variability. Complementarily, Figure 17 reports the MPJPE for each action averaged across all models and camera viewpoints. Together, these results indicate that the test set covers a broad and realistic distribution of actions and viewpoints (see also Figure 10), supporting the diversity required to evaluate generalization.